Unknowns for 1H NMR Final Report 4) Chemical Formula: CH360 IR: strong broad peak at 3300cm PPM 5) Chemical Formula: CsHN IR:If someone explain how to find these unknowns..And what they are and how you got the answer. I am having a hard time really understanding NMR and IR.. Thanks in advance
Answer #1

4) The formula is C4H10O

  • Degree of Unsaturation = (2*C+2-H)/2= (2*4+2-10)/2 = 0
  • This means that no double bonds are present in the molecule.
  • A compound with Oxygen having no double bond can either be ether or alcohol.
  • Further, the IR peak is at 3300 cm-1 . This suggests an alcohol group.
  • Further, in nmr spectroscopy too, the peak lying aroung 4.25 ppm suggests an alcohol group (-OH)
  • Further, a peak containing 6H at around 0.9 ppm suggests two methyl groups (-CH3).
  • A peak containing 2H at around 3.4 ppm suggests a -CH2- group. Further, the chemical shift is high which suggests an electronegative atom attached to the carbon (Oxygen). This means that -CH2OH must be present in the compound.
  • A peak containing 1H at around 1.4 ppm suggests a CH group. This peak is a multiplet which suggests adjoining carbons have multiple Hydrogens.
  • Putting together everything, we get the following structure-

I I-OOMI I OH I– All Hydrogens shown Line notation

The compound is 2-Methyl-1-propanol as shown in the figure. The two methyl group has NMR of 0.9 ppm at 6H. The CH group has 1H at 1.4 ppm. The adjoining CH2 group has 2H at 3.4 ppm and the Hydrogen of the alcohol group has nmr of 4.25 ppm